# Engagement-Driven Content Generation with Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) exhibit significant persuasion capabilities in one-on-one interactions, but their influence within social networks remains underexplored. This study investigates the potential social impact of LLMs in these environments, where interconnected users and complex opinion dynamics pose unique challenges. In particular, we address the following research question: can LLMs learn to generate meaningful content that maximizes user engagement on social networks?
To answer this question, we define a pipeline to guide the LLM-based content generation which employs reinforcement learning with simulated feedback. In our framework, the reward is based on an engagement model borrowed from the literature on opinion dynamics and information propagation. Moreover, we force the text generated by the LLM to be aligned with a given topic and to satisfy a minimum fluency requirement.
Using our framework, we analyze the capabilities and limitations of LLMs in tackling the given task, specifically considering the relative positions of the LLM as an agent within the social network and the distribution of opinions in the network on the given topic. Our findings show the full potential of LLMs in creating social engagement. Notable properties of our approach are that the learning procedure is adaptive to the opinion distribution of the underlying network and agnostic to the specifics of the engagement model, which is embedded as a plug-and-play component. In this regard, our approach can be easily refined for more complex engagement tasks and interventions in computational social science.
The code used for the experiments is publicly available at https://anonymous.4open.science/r/EDCG/.
Keywords Large Language Models  $\cdot$ Engagement Maximization  $\cdot$ Reinforcement Learning  $\cdot$ Opinion Dynamics
@@ 1 Introduction
Large Language Models (LLMs) have recently gained a great deal of attention, especially for their capabilities in one-to-one interactions. Researchers have deeply investigated their skills and limits, comparing LLMs to humans in one-on-one interactions such as, e.g., their abilities in gaming settings (Jacob et al., 2023; Fontana et al., 2024). A recent research direction empirically investigates the susceptibility (Chen et al., 2024) and persuasive capabilities of LLMs (Breum et al., 2024; Salvi et al., 2024), showing that contextual information in specific domains can enhance LLMs’ persuasiveness (Matz et al., 2024). In particular, LLMs have shown competence in generating persuasive messages in public health (Karinshak et al., 2023) and personalized advertising (Meguellati et al., 2024).
Figure 1: A visual representation of the proposed framework: (a) The process begins with a query $q$ prompted to the LLM, which generates content $t$. (b) The content $t$ is injected into the social network (i.e., posted by a node that corresponds to the LLM agent) and starts to propagate into the network according to an engagement model. (c) The number of active users at the end of the propagation is returned to the LLM as the observed reward. The process repeats until convergence.
@@
This body of work is limited to one-on-one interactions and the corresponding abilities of LLMs in isolated contexts. By contrast, the potential of LLMs within broader interconnected structures, such as their capability to influence social network users, is still largely unexplored. The ability to generate interactions that resonate across social networks can have significant implications, such as the potential to foster discussions, shape opinions, and drive behavioral changes among users. Following this observation, we tackle the following research question: can LLMs learn to generate meaningful content that maximizes user engagement on social networks?
@@
Our proposal to answer this question is a framework based on reinforcement learning, whose high-level vision is depicted in Figure 1. We start prompting to the LLM a query $q$ describing the topic of interest. The LLM generates a text $t$ that is then injected in the social network by an agent (a node of the network), which plays the role of the injection point and which we can think about as the LLM itself. The content $t$ then propagates through the network creating a certain level of engagement, i.e., the number of users that interacts with $t$. The observed level of engagement is then returned to the LLM as the observed reward and used to fine-tune the generation abilities of the LLM, aimed at maximizing the expected reward. This process is repeated till convergence.
@@
The framework presented in Figure 1 can be realized by means of Reinforcement Learning with Human Feedback (RLHF), an approach which has gained popularity due to its effectiveness in various learning tasks (Kaufmann et al., 2023), including fine-tuning of LLMs (Ouyang et al., 2024). However, conducting actual experiments involving content generation, propagation through a network, and waiting for responses is impractical due to the time and effort required. This is a common challenge in RLHF frameworks (Casper et al., 2024). To address this issue, we use a proxy model to simulate the network’s response, employing a formal engagement mechanism, which borrows ideas from both information propagation and opinion dynamics literature (Deffuant et al., 2000). In practice, we leverage an approach based on Reinforcement Learning from Simulated Feedback (RLSF), which enables efficient and effective feedback loop between the LLM agent and the network under study.
Thanks to the feedback loop, the LLM learns to adapt to the current topic, the social network structure, and the distribution of opinions in the network. For instance, if the opinion of the network on the given topic is prevalently negative, the generated content should have a negative leaning so to engage as many users as possible. Besides adherence to the given topic, we require a minimum fluency standard to the generated content.
@@
The rest of the paper is structured as follows. Section 2 reports an overview of the literature related to our problem. We formalize the problem in section 3 and provide the technical details of the approach in section 4. Section 5 reports the experiments aimed at evaluating the proposed approach, and finally in section 6 we discuss potential extensions and limitations.
@@ 2 Background and Related Work
LLM-Agents in complex environments. The integration of Agent-Based Models (ABMs) and Large Language Models has been an emerging area of research. Notable studies, such as those by Vezhnevets et al. (2023) and Gurcan (2024), focus on utilizing LLMs as social agents within ABMs to simulate micro-level behaviors and observe macro-level patterns. These micro-level behaviors can correspond to opinion dynamics models (Chuang et al., 2023), and the resulting macro-level patterns can lead to known universal laws such as scale-free networks (De Marzo et al., 2023). Recent studies have started to place LLM agents into more complex environments to assess their ability to mimic human behavior (Park et al., 2023), and to understand their capabilities in negotiation (Bianchi et al., 2024), code development (Qian et al., 2023), and even in tasks like building houses (Chen et al., 2023). A concurrent line of inquiry is whether LLMs can learn to align their behavior with social norms (Fränken et al., 2023).
@@
Social media environments. The adoption of information diffusion models in data mining methodologies has been thoroughly investigated for designing viral marketing strategies (Kempe et al., 2003; Leskovec et al., 2007; Chen et al., 2010; Lu et al., 2013; Barbieri et al., 2013; Barbieri and Bonchi, 2014; Tu and Neumann, 2022). Various optimization problems have been tackled such as selecting influential seeds (Kempe et al., 2003; Tu and Neumann, 2022), designing product features (Barbieri and Bonchi, 2014), selecting seeds for fair competition (Lu et al., 2013), or seed selection with minimum regret (Aslay et al., 2015). All these works highlight the importance of considering the underlying network structure in the spreading process - a concept that out-of-the-box LLMs are not equipped to consider. Most of these optimization problems often do not account for the fact that social media operates with real texts. Therefore, the fluency and realism of the content need to be modeled and taken into consideration, generating text that inherently possesses viral characteristics (as described in Aral and Walker (2011)).
Proxy for Human Feedback. Regardless of the type of objectives that machine learning models have, a correct reward function is necessary. In particular, the challenges associated with human feedback in evaluating LLMs have led to alternative approaches, such as LLM-as-a-Judge (Huang et al., 2024; Verga et al., 2024) and human-in-the-loop methods (Boubdir et al., 2023; Amirizaniani et al., 2024), which aim to enhance scalability and efficiency. While our work shares the goal of reducing the reliance on human experts for data annotation, it introduces a fundamentally novel approach by simulating context-dependent feedback that incorporates the social network structure. Hence, this innovation also minimizes the need for slow, costly, and potentially unethical online experiments, offering a more efficient and ethically sound alternative. To the best of our knowledge, no existing work leverages simulated social media engagement models as feedback mechanisms to fine-tune Large Language Models, addressing a critical gap in current research.
@@ 3 Engagement Model
We are given a social network represented as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where nodes are users and a directed edge $(u,v)\in\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ indicates that $v$ is a “follower” of $u$: thus $u$ can propagate content to $v$, but not necessarily vice versa. We are also given a vector $\vec{x}\in[0,1]^{|\mathcal{V}|}$, associating each node $u\in\mathcal{V}$ with a scalar $x_{u}$ between 0 and 1, representing its opinion on a given topic. We interpret $x_{u}\approx 1$ as being in favor of the topic, while $x_{u}\approx 0$ indicates disapproval of the topic.

### v4

@@ Abstract
Large Language Models (LLMs) demonstrate significant persuasive capabilities in one-on-one interactions, but their influence within social networks, where interconnected users and complex opinion dynamics pose unique challenges, remains underexplored. This paper addresses the research question: Can LLMs generate meaningful content that maximizes user engagement on social networks?
To answer this, we propose a pipeline using reinforcement learning with simulated feedback, where the network’s response to LLM-generated content (i.e., the reward) is simulated through a formal engagement model. This approach bypasses the temporal cost and complexity of live experiments, enabling an efficient feedback loop between the LLM and the network under study. It also allows to control over endogenous factors such as the LLM’s position within the social network and the distribution of opinions on a given topic. Our approach is adaptive to the opinion distribution of the underlying network and agnostic to the specifics of the engagement model, which is embedded as a plug-and-play component. Such flexibility makes it suitable for more complex engagement tasks and interventions in computational social science.
Using our framework, we analyze the performance of LLMs in generating social engagement under different conditions, showcasing their full potential in this task. The experimental code is publicly available. [^1] https://github.com/mminici/Engagement-Driven-Content-Generation
Keywords Large Language Models  $\cdot$ Engagement Maximization  $\cdot$ Reinforcement Learning
@@ 1 Introduction
Large Language Models (LLMs) have recently gained a great deal of attention, especially for their capabilities in one-to-one interactions. Researchers have deeply investigated their skills and limits, comparing LLMs to humans in one-on-one interactions such as, e.g., their abilities in gaming settings (Jacob et al., 2024; Fontana et al., 2024). A recent research direction empirically investigates the susceptibility (Chen et al., 2024) and persuasive capabilities of LLMs (Breum et al., 2024; Salvi et al., 2024), showing that contextual information in specific domains can enhance LLMs’ persuasiveness (Matz et al., 2024). In particular, LLMs have shown competence in generating persuasive messages in public health (Karinshak et al., 2023) and personalized advertising (Meguellati et al., 2024).
@@
This body of work focuses on the capabilities of LLMs in one-on-one interactions within isolated contexts. However, their potential within broader interconnected structures, such as their ability to create information cascades among the users of a social network, remains largely unexplored. Such capacity can have significant implications, including fostering discussions, shaping opinions, and driving behavioral changes among users. In light of this, we address the following research question: Can LLMs learn to generate meaningful content that maximizes user engagement on social networks?
@@
Our proposal to answer this question is a framework based on reinforcement learning, whose high-level vision is depicted in Figure 1. We begin by prompting the LLM with a query $q$ describing the topic of interest. The LLM generates a text $t$, which is then injected into the social network by an agent (a node of the network), acting as the injection point and representing the LLM itself. The content $t$ propagates through the network, generating a certain level of engagement, i.e., the number of users interacting with $t$. This observed engagement, combined with a fluency score of the text $t$, is returned to the LLM as a reward which is used to fine-tune the LLM’s generation abilities to maximize the expected reward. This process is repeated until convergence.
@@
The framework presented in Figure 1 can be realized by means of Reinforcement Learning with Human Feedback (RLHF), an approach which has gained popularity due to its effectiveness in various learning tasks (Kaufmann et al., 2023), including fine-tuning of LLMs (Ouyang et al., 2024). However, conducting actual live experiments involving propagating content through a social network and waiting for responses, is complex and time consuming. This is a common challenge in RLHF frameworks (Casper et al., 2024). To overcome this, we leverage an approach based on Reinforcement Learning from Simulated Feedback (RLSF). Specifically, we adopt a proxy model to simulate the network’s response, employing a formal engagement mechanism, which borrows ideas from both information propagation and opinion dynamics literature (Deffuant et al., 2000). In addition to bypassing the temporal cost and complexity of live experiments, thus enabling an efficient and effective feedback loop between the LLM agent and the network under study, such an approach also allows to control over endogenous factors such as the LLM’s position within the social network and the distribution of opinions on a given topic. Thanks to the feedback loop, the LLM learns to adapt to the current topic, the social network structure, and the distribution of opinions in the network. For instance, if the opinion of the network on the given topic is prevalently negative, the generated content should have a negative leaning so to engage as many users as possible.
Figure 1: A visual representation of the proposed framework: (a) A query $q$ is prompted to the LLM, which generates content $t$. (b) The content $t$ is injected into the social network (i.e., posted by a node that corresponds to the LLM agent) and starts to propagate into the network according to an engagement model. (c) The number of active users at the end of the propagation is returned to the LLM as the observed reward. The process repeats until convergence.
@@
The rest of the paper is structured as follows. Section 2 reports an overview of the literature related to our problem. We formalize the problem in Section 3 and provide the technical details of the approach in Section 4. Section 5 and 6 outline the experimental design used to evaluate the proposed methodology and analyze how the experimental results address the formulated research questions. Finally, in Section 7, we discuss the implications of our findings, along with potential extensions and limitations of the study.
@@ 2 Background and Related Work
LLM-Agents in complex environments. The integration of Agent-Based Models (ABMs) and Large Language Models has been an emerging area of research. Notable studies, such as those by Vezhnevets et al. (2023) and Gürcan (2024), focus on utilizing LLMs as social agents within ABMs to simulate micro-level behaviors and observe macro-level patterns. These micro-level behaviors can correspond to opinion dynamics models Chuang et al. (2024), and the resulting macro-level patterns can lead to known universal laws such as scale-free networks De Marzo et al. (2023). Recent studies have started to place LLM agents into more complex environments to assess their ability to mimic human behavior Park et al. (2023), and to understand their capabilities in negotiation Bianchi et al. ([n. d.]), code development Qian et al. (2024), and even in tasks like building houses Chen et al. (2023). A concurrent line of inquiry is whether LLMs can learn to align their behavior with social norms Fränken et al. ([n. d.]).
@@
Social media environments. The adoption of information diffusion models in data mining methodologies has been thoroughly investigated for designing viral marketing strategies Kempe et al. (2003); Leskovec et al. (2007); Chen et al. (2010); Lu et al. (2013); Barbieri et al. (2013); Barbieri and Bonchi (2014); Tu and Neumann (2022). Various optimization problems have been tackled such as selecting influential seeds Kempe et al. (2003); Tu and Neumann (2022), designing product features Barbieri and Bonchi (2014), or other variations Lu et al. (2013); Aslay et al. (2015). Yet, a crucial practical dimension remains largely unexamined: how to compose textual content that effectively propagates through a network. In this work, we bridge this gap by proposing a methodology that enables LLMs to generate content maximizing engagement on social platforms.
Proxy for human feedback. Regardless of the type of objectives that machine learning models have, a correct reward function is necessary. In particular, the challenges associated with human feedback in evaluating LLMs have led to alternative approaches, such as LLM-as-a-Judge Huang et al. (2024); Verga et al. (2024) and human-in-the-loop methods Boubdir et al. (2023); Amirizaniani et al. (2024), which aim to enhance scalability and efficiency. While our work shares the goal of reducing the reliance on human experts for data annotation, it introduces a fundamentally novel approach by simulating context-dependent feedback that incorporates the social network structure. Hence, this innovation also minimizes the need for slow, costly, and potentially unethical online experiments, offering a more efficient and ethically sound alternative. To the best of our knowledge, no existing work leverages simulated social media engagement models as feedback mechanisms to fine-tune LLMs, addressing a critical gap in current research.
@@ 3 Engagement Model
As discussed in Section 1, we adopt a Reinforcement Learning from Simulated Feedback approach. This method not only helps us avoiding the complexity of live experiments but also allows us to control contextual settings, such as the distribution of opinions and the LLM’s position within the social network. Given our goal to study the capabilities of LLMs in generating content that resonates across the network, we propose a novel engagement model which borrows from information propagation Kempe et al. (2003) the cascade mechanism, and from opinion dynamics Deffuant et al. (2000), the single engagement mechanism.
@@
More formally, we consider a social network represented as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where nodes are users and a directed edge $(u,v)\in\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ indicates that $v$ is a “follower” of $u$: thus $u$ can propagate content to $v$, but not necessarily vice versa. We are also given a vector $\vec{x}\in[0,1]^{|\mathcal{V}|}$, associating each node $u\in\mathcal{V}$ with a scalar $x_{u}$ between 0 and 1, representing its opinion on a given topic. We interpret $x_{u}\approx 1$ as being in favor of the topic, while $x_{u}\approx 0$ indicates disapproval of the topic.